OpenStudy (mstv1112):

1 questions Will give MEDAL!

OpenStudy (anonymous):

what

OpenStudy (mstv1112):

Show work please. Sorry it's sloppy. lol.

OpenStudy (mstv1112):

\[3(10 - 6)^{2}\]

OpenStudy (alexandervonhumboldt2):

3*(100-2*10*6+36)=

OpenStudy (alexandervonhumboldt2):

can you do adding and substaraction?

OpenStudy (alexandervonhumboldt2):

48

OpenStudy (alexandervonhumboldt2):

medal fan please

OpenStudy (anonymous):

48

OpenStudy (mstv1112):

Im sorry I don't understand how you got that can you show me steps? I'm just learning this...

OpenStudy (alexandervonhumboldt2):

ok (a-b)^2=a^2-2ab+b^2

OpenStudy (alexandervonhumboldt2):

got it?

OpenStudy (mstv1112):

Kinda ....

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

me want medal